OpenStudy (anonymous):

3 5/7 = 35/7 105/7 38/7 26,7

OpenStudy (anonymous):

umm what are you asking exactly?

OpenStudy (anonymous):

???

OpenStudy (kymber):

What's the question?

OpenStudy (anonymous):

another way of 3 5/7

OpenStudy (anonymous):

3 x 7 = 21 + 5 = 26/7

OpenStudy (anonymous):

kymber is that right?

OpenStudy (anonymous):

ok if you get 7/7 that = 1 So 3 is 21/7 and 21 + 5 = 26/7

OpenStudy (kymber):